OpenStudy (anonymous):

write a formula for the apparent nth term of the sequence ( just 2 more of these types) 1, -1/2, 1/3, -1/4, 1/5 (the other one) 4, 2, 4/3, 1, 4/5

OpenStudy (callisto):

@lakisha2013 What have you got for this question?

OpenStudy (anonymous):

i have been questions but they are hard i have not got an answer the last time i did this was about 2 years ago or more

OpenStudy (callisto):

I meant what have you done for this question... It doesn't matter you get the answer or not, but at least show us what you've tried. It's easier to help when we know what you don't understand.

OpenStudy (anonymous):

i have been trying to find the patterns but i cannot seem to find them

OpenStudy (callisto):

1 = [(-1)^(1-1)] (1/1) -1/2 = [(-1)^(2-1)] (1/2) 1/3 = [(-1)^(3-1)] (1/2) -1/4 = ...? 1/5 =...? Can you work it out now?

OpenStudy (anonymous):

no because when i did the third problem i did not get 1/3

OpenStudy (callisto):

Sorry, it's 1/3 = [(-1)^(3-1)] (1/3)

OpenStudy (anonymous):

so the next one would be [(-1)^(4-1)](1/4)

OpenStudy (callisto):

Yes.

OpenStudy (callisto):

Now, can you write it in general terms?

OpenStudy (anonymous):

it would be( -1^(n-1))(1/n)

OpenStudy (callisto):

Yes!

OpenStudy (anonymous):

i have started on the other one i got that the first one is either subtracting 2 or dividing 2

OpenStudy (callisto):

So, for the other one, 4 = 4/1 2 = 4/2 4/3 = 4/3 1 = ... ? 4/5 = ...?

OpenStudy (anonymous):

it would be 4/n

OpenStudy (callisto):

Yes! Sorry.. I couldn't see your post. For the second question, it can hardly involve subtraction because of the denominator... (it varies!)

OpenStudy (anonymous):

ok thank

OpenStudy (callisto):

Welcome!